# Thermal condensation with polybaric pressure

### v1

@@ 4 The linear growth phase
Figure 1: Linear evolution of condensation and wave modes with their growth rates estimated from simulation. (a) both modes unstable (top left), (b) only condensation mode unstable (right). The corresponding growth rates $\varpi_{c,w}$ are shown in the figure, as determined from the simulation. The blue colored graphs (solid and circle) indicate data from simulation and the red colored solid lines indicate an exponential fitting corresponding to the linear growth phase.
@@
$k^{2}K+\mathfrak{L}_{T}+\frac{1}{\gamma-1}\mathfrak{L}_{\rho}$ $<$ $0$ (60)
@@
(a) both $\begin{array}[]{lcll}\textrm{(a) both unstable}&:&\mathfrak{L}_{T}<0,|A|<\mathfrak{L}_{\rho}<{\displaystyle\frac{2}{3}}|A|,&A=k^{2}K+\mathfrak{L}_{T},\\ \textrm{(b) only wave}&:&\mathfrak{L}_{\rho}<0,-B<\mathfrak{L}_{T}<-B-{\displaystyle\frac{5}{2}}\mathfrak{L}_{\rho},&B=k^{2}K-\mathfrak{L}_{\rho},\\ \textrm{(c) only condensation}&:&\mathfrak{L}_{\rho}>A,&A>0,\\ &&\mathfrak{L}_{\rho}>{\displaystyle\frac{2}{3}}|A|,&A<0.\end{array}$ (61)
In Fig.1, we have shown the evolution of a single wave in the 1D case for two cases– both modes unstable and only condensation mode unstable, along with the corresponding numerically estimated growth rates $\varpi_{c,w}$ of the condensation and the wave modes. In all the panels of Fig.1, the blue colored curves (solid and circles) denote the results from simulation and the red colored solid line shows an exponential fitting for the linear growth phase. The left panel of Fig.1 shows the evolution of a single wave when both wave and condensation modes are unstable. While the primary figure shows the evolution of the condensation mode with a growth rate of $\varpi_{c}\sim 3.57\times 10^{-3}$ (normalized), the inset shows the evolution of the wave mode with a growth rate of $\varpi_{w}\sim 1.52\times 10^{-3}$, determined from the actual growth of amplitude of the wave after removal of the condensation growth. The right panel of Fig.1 shows the evolution of the condensation mode with a growth rate of $\varpi_{c}\sim 5.29\times 10^{-3}$. The inset shows a zoomed out portion of the initial evolution, where one can clearly see a stable wave mode with only the condensation mode growing. The evolution of the wave mode, corresponding to (b) in Eq.(61) is found to have a growth rate $\varpi_{w}\sim 1.3\times 10^{-3}$ (not shown in figure). The analytically calculated corresponding growth rates $\left(\varpi_{c,w}^{{\rm anal}}\right)$ for these cases are respectively (a) $\varpi_{c}^{{\rm anal}}\sim 3.56\times 10^{-3}$, $\varpi_{w}^{{\rm anal}}\sim 1.52\times 10^{-3}$, (b) $\varpi_{w}^{{\rm anal}}\sim 1.31\times 10^{-3}$, (c) $\varpi_{c}^{{\rm anal}}\sim 5.27\times 10^{-3}$. As we can see, the growth rates determined from the simulation are almost in complete agreement with the analytically calculated ones.
@@ 5 Nonlinear polybaric evolution
In this section, we are going to present our primary simulation results with polybaric pressure under a constant magnetic field. All the simulations are carried out in a 2D geometry with a simulation box size of $(20\pi\times 20\pi)\,{\rm pc}^{2}$. The simulation is started with random initial perturbation in the density map, which is slightly unstable to the thermal condensation mode. The perturbation amplitude is set at $\sim 1\%$ of the equilibrium density.
@@ 5.1 No magnetic field
In Fig.2, we have shown the evolution of the density map with scalar pressure from a random initial perturbation, which is unstable to the thermal condensation mode. The plots show the thermal condensation in density resulting in formation of cold neutral matter (CNM). The small rectangular box in each panel indicates the area, which is used to calculate the initial (linear) growth rate of the instability. The linear growth of the TI is shown in the right panel of the bottom row of Fig.2. The blue circles indicate the density evolution (within the rectangular box) and the red curve is an exponential fit corresponding to the linear regime. For the ISM parameters already mentioned before and the equilibrium state that we have used, the maximum linear growth rate (with respect to perturbation wave number) that is possible, comes out to be $\varpi_{{\rm max}}^{{\rm theory}}\sim 1.03\times 10^{-2}\,{\rm Myr^{-1}}$, which we are expected to observe during a simulation with random perturbation. As shown in Fig.2 (right panel, bottom row), the linear growth rate in this case is found to be $\varpi^{{\rm simul}}\sim 1.93\times 10^{-2}\,{\rm Myr^{-1}}$, which closely agrees with the theoretical value. The density $\rho$ then increases from an equilibrium value of unity (normalized to the equilibrium value $\rho_{0}$) to about an average value of $\sim 180$ during the nonlinear phase.
@@
Figure 2: The evolution of the density map under TI. The timestamp in different simulation panels are inscribed within each panel. The initial state of the simulation is a random density perturbation which is unstable to the thermal condensation mode. The small rectangle shown in each of the top four panels indicates a density condensation area, which is used to calculate the growth rate $(\varpi)$, shown in the right panel of the bottom row. The left panel of the bottom row shows temperature perturbations corresponding to density perturbations at $200\,{\rm Myr}$.
@@
As one can see that the nonlinear evolution starts at about $\sim 500\,{\rm Myr}$ from a very small value (just $\sim 1\%$of the equilibrium density). Naturally, the instability will only linearly grow in the initial phase. We should note that if a higher perturbation amplitude is introduced at the initial stage, the nonlinear evolution will occur sooner. As such, the onset of nonlinear evolution and hence the time required to attain the final condensation stage is quite arbitrary and may depend on the particular simulation concerned. What is however invariant is the linear growth rate of the instability $\varpi$, which is quite consistent with the analytical estimate, as already shown in the previous section and in the beginning of this section. We also note that the nonlinear evolution never attains a completely steady state and fluctuates around a mean value and hence the size and structure of the CNMs will fluctuate with time but the overall density structures have a nonlinear evolution timescale of $\sim 10\,{\rm Myr}$. Note that the density shown in this plot is a cumulative density within that rectangle. In left panel of the bottom row of Fig.2, we have also shown the normalized temperature map $(T/T_{0})$ at a timestamp of $200\,{\rm Myr}$, which clearly shows the anti-correlation between the corresponding density map and formation of CNM and warm neutral matter (WNM), which further confirms the process of thermal condensation with pressure equilibration.
5.2 Constant magnetic field
@@ 5.2.1 Plasma $\beta$
Figure 3: Density map for $a_{p}=0$ and $\beta=10^{4}$ (‘weak’ field) for different phases of the TI (top four panels). While the top three panels show the evolution for $a_{p}=0$, the right panel of the second row shows the same for $a_{p}=3$. The bottom row shows the evolution at the end of $850\,{\rm Myr}$ for $a_{p}=0$ and $3$ for the case of ‘strong’ field $(\beta=5)$.
   Refer to caption          Refer to caption
Figure 4: The velocity field for $\bm{B}=0$ (top row) and ‘strong’ field (second row) with $\beta=5$ and $a_{p}=3$ at different time stamps. The local Mach number of the plasma flow for the case $\beta=5$ at $850\,{\rm Myr}$ is shown in the left panel of the third row. The bottom two panels on the right show the power spectrum density (PSD) of the velocity field for $\bm{B}>0\,(\beta=5)$ (top) and $\bm{B}=0$ (bottom). The respective scalings with the spatial wave number are inscribed in the panels. The left bottom panel shows the compensated PSDs of the two cases. The yellow and black colored spectra in both panels indicate the solenoidal and compressible parts that comprise the total spectra.
@@ 5.2.2 Thermal condensation
In Fig.3, we have shown the evolution of the density map for the ‘weak’ field $(\beta=10^{4})$ for the parameter $a_{p}=0$ and $3$ (top four panels). The initial state is the same as before (as in the case of scalar pressure). What we can see from the figure is that the initial evolution of density is predominantly along the ambient magnetic field and the final condensations are also developed parallel to the magnetic field. However, as the evolution becomes fully nonlinear, the condensations become diffused compared to the similar map for scalar pressure (see Fig.2) though the final condensations seem to to be aligned with the ambient magnetic field. While all the top four panels in Fig.3 refer to the ‘weak’ field case for $a_{p}=0$ and $3$, the lowermost panel of Fig.3 shows the evolution of density after $850\,{\rm Myr}$ for $a_{p}=0$ and $3$ for the ‘strong’ field case. We can see that for comparatively higher $p_{\parallel}$ (larger $a_{p}$), we have slightly lower condensations.
@@
We now analyse what exactly happens in the case of a strong field. In particular, we choose the case for $\beta=5$ and $a_{p}=3$. We have chosen to plot the velocity field vectors in Fig.4 which is superimposed on a velocity field map to highlight the plasma flow. Note that in Fig.4, the lengths of the velocity vectors are kept uniform for clarity, while the background velocity field indicates the strength of the net plasma flow at that point. The top row of Fig.4 shows the velocity field vectors and map for scalar pressure when there is no ambient magnetic field. We can see that in this case, the plasma velocity is more or less ordered, driven by the density condensations. During the fully developed nonlinear stage of the evolution at $t\sim 850\,{\rm Myr}$, the bulk plasma flow is quite ordered except at the site of the condensations. As we can see that the density condensations are also closely followed by higher plasma flow into the regions of higher mass densities. In contrast to this, when there is a magnetic field present, initially most of the plasma flow is along the magnetic field lines. The low density areas are marked by small-magnitude perpendicular plasma flow, as expected. This also causes initial density striations to form along the magnetic field lines. We note that the immediate effect of a magnetic field is to restrict the plasma flow in the transverse direction, which points to the fact that condensation always occurs along the magnetic field lines as the parallel velocity is mostly unrestricted. However, even a small transverse velocity component can make a considerable difference in the final condensation state (Hennebelle & Pérault, 1999). As a result, we can expect that the thermal instability will be stifled due to the flow restrictions imposed by the magnetic field. However, as the total energy content must be conserved, we certainly expect a field aligned, more diffused condensation state and the resultant velocity field will not be highly ordered leading to a turbulent-like plasma flow, as can be seen from the right figure of the second row of Fig.4 (see next subsection).

### v2

@@ 4 The linear growth phase
Figure 2: Top left panel (a) shows normalized growth rates of the condensation mode for different values of conductivity coefficient (points from simulation are shown by blue circles). Rest of the panels show linear evolution of condensation and wave modes with their growth rates estimated from simulation. (a) both modes unstable [panel (b)], (b) only wave mode unstable [panel (c)], (c) only condensation mode unstable [panel (d)]. The corresponding growth rates $\varpi_{c,w}$ are shown in the figure, as determined from the simulation. The blue colored graphs (solid and circle) indicate data from simulation and the red colored solid lines indicate an exponential fitting corresponding to the linear growth phase.
@@
$k^{2}K+\mathfrak{L}_{T}+\frac{1}{\gamma-1}\mathfrak{L}_{\rho}$ $<$ $0.$ (63)
@@
$\begin{array}[]{lcll}\textrm{(\emph{a}) both unstable}&:&\mathfrak{L}_{T}<0,|A|<\mathfrak{L}_{\rho}<{\displaystyle\frac{2}{3}}|A|,\\ &&A=k^{2}K+\mathfrak{L}_{T},\\ \textrm{(\emph{b}) only wave}&:&\mathfrak{L}_{\rho}<0,\\ &&-A_{1}<\mathfrak{L}_{T}<-A_{1}-{\displaystyle\frac{5}{2}}\mathfrak{L}_{\rho},\\ &&A_{1}=k^{2}K-\mathfrak{L}_{\rho},\\ \textrm{(\emph{c}) only condensation}&:&\mathfrak{L}_{\rho}>A,A>0,\\ &&\mathfrak{L}_{\rho}>{\displaystyle\frac{2}{3}}|A|,A<0.\end{array}$ (64)
To benchmark our numerical scheme, we use the linear dispersion relation
$\omega^{3}-\gamma\omega k^{2}+i(\gamma-1)\left[\omega^{2}\left(k^{2}K+\mathfrak{L}_{T}\right)\right$.
$\left.-k^{2}\left(k^{2}K+\mathfrak{L}_{T}-\mathfrak{L}_{\rho}\right)\right]$ $=$ $0$ (65)
by considering a perturbation of the form $\sim e^{-i(\omega t-kx)}$.
In Fig.2, we compare the theoretical growth rates calculated using the linear dispersion relation with the growth rate measured from 1D TI simulation. In the same figure, the evolution of a single wave for all the above three conditions along with the corresponding numerically estimated growth rates $\varpi_{c,w}$ of the condensation and the wave modes are also shown. In all the panels of Fig.2, the blue colored curves (solid and circles) denote the results from simulation and the red colored solid line shows an exponential fitting for the linear growth phase. In Fig.2(a), normalized growth rates of the condensation mode calculated from the linear dispersion equation for different values of conductivity coefficient $(K=0,0.01,0.1\,\rm{and}\,1)$ are shown along with growth rates calculated from simulation for $K=0.01$. Fig.2(b) shows the evolution of a single wave when both wave and condensation modes are unstable. While the primary figure shows the evolution of the condensation mode with a growth rate of $\varpi_{c}\sim 3.57\times 10^{-3}$ (normalized), the inset shows the evolution of the wave mode with a growth rate of $\varpi_{w}\sim 1.52\times 10^{-3}$, determined from the actual growth of amplitude of the wave after removal of the condensation growth. Fig.2(c) shows the evolution of the wave mode (displayed as a difference between the maximum and minimum amplitudes of the wave over time) with a growth rate $\varpi_{w}\sim 1.3\times 10^{-3}$. Finally, Fig.2(d) shows the evolution of the condensation mode with a growth rate of $\varpi_{c}\sim 5.29\times 10^{-3}$. The inset shows a zoomed in portion of the initial evolution, where one can clearly see a stable wave mode with only the condensation mode growing. The analytically calculated growth rates $\left(\varpi_{c,w}^{{\rm an}}\right)$ for the three cases from Eq.(65) are respectively (a) $\varpi_{c}^{{\rm an}}\sim 3.56\times 10^{-3}$, $\varpi_{w}^{{\rm an}}\sim 1.52\times 10^{-3}$, (b) $\varpi_{w}^{{\rm an}}\sim 1.31\times 10^{-3}$, (c) $\varpi_{c}^{{\rm an}}\sim 5.27\times 10^{-3}$. As we can see, the growth rates determined from the simulation are almost in complete agreement with the analytically calculated ones.
@@ 5 Nonlinear polybaric evolution
In this section, we are going to present our primary simulation results with polybaric pressure under a constant magnetic field. All the simulations are carried out in a 2D geometry with a simulation box size of $(20\pi\times 20\pi)\,{\rm pc}^{2}$. We use the supernova forcing term as a source of perturbation which is fed into the simulation at the initial stage $t_{0}=0$ through a randomised tunable velocity field with pre-determined ratio of solenoidal and compressive components.
5.1 Supernova-driven velocity field
We note that supernovæ explosions create a radial flow with an expanding shock into the immediate vicinity. However, as time progresses, at larger scale these expanding shocks begin to create vorticity and generate turbulence which becomes akin to random fluctuations, which is also supported by MHD simulations (Korpi et al., 1999). So, we assume that the supernova forcing can be approximated as random fluctuation so far as the large-scale ISM structure is concerned (Korpi et al., 1999; Padoan et al., 2016).
So, in order to model the supernova-driven energy injection, we use a velocity field with a mix of solenoidal and compressive parts (Padoan et al., 2016). In particular, we use three scenarios — Case (i): a velocity field with a relatively dominant compressive part $(\sim 54\%)$, which can be thought to be like a remnant of supernova shock-driven field, Case (ii): one with a dominant solenoidal field $(\sim 72\%)$ and weak compressive part $(\sim 28\%)$, and Case (iii): one with a solenoidal part $(\gtrsim 95\%)$ only with a negligible amount of compressive part. We shall observe that a compressive velocity field helps drive the TI relatively quicker in reaching a nonlinear saturation compared to one with a dominant solenoidal part (Federrath et al., 2010), though statistically all velocity fields result in similar volume and mass fractions of multiphase ISM components, as shown in Subsection 5.3.5.
In Fig.3, we have shown the decomposed power spectrum density (PSD) of the net initial velocity for case (i). The velocity field has been decomposed through a Helmholtz-Hodge decomposition (Bhatia et al., 2013)
$\bm{v}=-\nabla\phi+\nabla\times A,$ (66)
where the $\nabla\phi$ is the compressive part of the velocity field and the other solenoidal. In what follows, we shall analyse our results for this case and in Section 5.3.5, we shall consider the other two cases.
Figure 3: The power spectrum density (PSD) of the initial velocity field used for modelling the supernova forcing term. This velocity field has a relatively dominant compressive part.
@@ 5.2 No magnetic field
In Fig.4, we have shown the evolution of the density map with scalar pressure from a random velocity field shown in case (i) above. The plots show the thermal condensation in density resulting in formation of cold neutral matter (CNM). The small rectangular box in each panel indicates the area, which is used to calculate the initial (linear) growth rate of the instability. The linear growth of the TI is shown in Fig.4(f). The blue circles indicate the density evolution (within the rectangular box) and the red curve is an exponential fit corresponding to the linear regime. For the ISM parameters already mentioned before and the equilibrium state that we have used, the maximum linear growth rate (with respect to perturbation wave number) that is possible, comes out to be $\varpi_{{\rm max}}^{{\rm theory}}\sim 1.03\times 10^{-2}\,{\rm Myr^{-1}}$, which we are expected to observe during a simulation with random perturbation. As shown in Fig.4(f), the linear growth rate in this case is found to be $\varpi^{{\rm simul}}\sim 1.93\times 10^{-2}\,{\rm Myr^{-1}}$, which closely agrees with the theoretical value. The density $\rho$ then increases from an equilibrium value of unity (normalized to the equilibrium value $\rho_{0}$) to about an average value of $\sim 180$ during the nonlinear phase.
@@
Figure 4: The evolution of the density map under TI. The timestamp in different simulation panels are inscribed within each panel. The initial state of the simulation is a random density perturbation, resulting out of supernovæ forcing (see text), which is unstable to the thermal condensation mode. The small rectangle shown in each of the top four panels indicates a density condensation area, which is used to calculate the growth rate $(\varpi)$, shown in panel (f). Panel (e) shows temperature perturbations corresponding to density perturbations at $200\,{\rm Myr}$ in panel (c).
@@
As one can see, the nonlinear evolution starts at about $\sim 500\,{\rm Myr}$ from a very small value. Naturally, the instability will only linearly grow in the initial phase. We should note that if a higher perturbation amplitude is introduced at the initial stage, the nonlinear evolution will occur sooner. As such, the onset of nonlinear evolution and hence the time required to attain the final condensation stage is quite arbitrary and may depend on the particular simulation concerned. What is however invariant is the linear growth rate of the instability $\varpi$, which is quite consistent with the analytical estimate, as already shown in the previous section and in the beginning of this section. We also note that the nonlinear evolution never attains a completely steady state and fluctuates around a mean value and hence the size and structure of the CNMs will fluctuate with time but the overall density structures have a nonlinear evolution timescale of $\sim 10\,{\rm Myr}$. Note that the density shown in this plot is a cumulative density within that rectangle. In Fig.4(e), we have also shown the normalized temperature map $(T/T_{0})$ at a timestamp of $200\,{\rm Myr}$, which clearly shows the anti-correlation between the corresponding density map and formation of CNM and warm neutral matter (WNM), which further confirms the process of thermal condensation with pressure equilibration.
5.3 Constant magnetic field and pressure anisotropy
@@ 5.3.1 Plasma $\beta$
Figure 5: Density map for $a_{p}=0$ and $\beta=10^{4}$ (‘weak’ field) for different phases of the TI [panels (a)–(c)]. While the first three panels show the evolution for $a_{p}=0$, panel (d) shows the same for $a_{p}=3$. The bottom row [panels (e) and (f)] shows the evolution at the end of $850\,{\rm Myr}$ for $a_{p}=0$ and $3$ for the case of ‘strong’ field $(\beta=5)$.
Figure 6: Density map for $\beta=0.1$ (‘very strong’ field) in the presence of thermal conduction with $a_{p}=3$ [panels (a) and (b)] and in the absence of thermal conduction with different values of pressure anisotropy parameter ($a_{p}=0,1,\,\rm{and}\,3$) [panels (c)–(h)]. The Right panels show a zoomed in view of the density maps in the corresponding left panels.
Figure 7: The velocity field for $\bm{B}=0$ [panels (a), (b)] and ‘strong’ field with $\beta=5$ and $a_{p}=3$ [panels (c), (d)] at different time stamps. The local Mach number of the plasma flow for the case $\beta=5$ at $850\,{\rm Myr}$ is shown in panel (e). The bottom two panels on the right show the power spectrum density (PSD) of the velocity field for $\bm{B}>0\,(\beta=5)$ [panel (f)] and $\bm{B}=0$ [panel (h)]. The respective scalings with the spatial wave number are inscribed in the panels. Panel (g) shows the compensated PSDs of the two cases. The yellow and black colored spectra in both panels indicate the solenoidal and compressive parts that comprise the total spectra.
@@ 5.3.2 Thermal condensation
In Fig.5, we have shown the evolution of the density map for the ‘weak’ field $(\beta=10^{4})$ for the parameter $a_{p}=0$ and $3$ (top four panels). The initial state is the same as before (as in the case of scalar pressure). What we can see from the figure is that the initial evolution of density is predominantly along the ambient magnetic field and the final condensations are also developed parallel to the magnetic field. However, as the evolution becomes fully nonlinear, the condensations become diffused compared to the similar map for scalar pressure (see Fig.4) though the final condensations seem to be aligned with the ambient magnetic field. While all the top four panels in Fig.5 refer to the ‘weak’ field case for $a_{p}=0$ and $3$, the lowermost panel of Fig.5 shows the evolution of density after $850\,{\rm Myr}$ for $a_{p}=0$ and $3$ for the ‘strong’ field case. We can see that for comparatively higher $p_{\parallel}$ (larger $a_{p}$), we have slightly lower condensations.
@@
We now analyse what exactly happens in the case of a strong field. In particular, we choose the case for $\beta=5$ and $a_{p}=3$. We have chosen to plot the velocity field vectors in Fig.7 which is superimposed on a velocity field map to highlight the plasma flow. Note that in Fig.7, the lengths of the velocity vectors are kept uniform for clarity, while the background velocity field indicates the strength of the net plasma flow at that point. The top row of Fig.7 shows the velocity field vectors and map for scalar pressure when there is no ambient magnetic field. We can see that in this case, the plasma velocity is more or less ordered, driven by the density condensations. During the fully developed nonlinear stage of the evolution at $t\sim 850\,{\rm Myr}$, the bulk plasma flow is quite ordered except at the site of the condensations. As we can see that the density condensations are also closely followed by higher plasma flow into the regions of higher mass densities. In contrast to this, when there is a magnetic field present, initially most of the plasma flow is along the magnetic field lines. The low density areas are marked by small-magnitude plasma flow, as expected. This also causes initial density striations to form along the magnetic field lines. We note that the immediate effect of a magnetic field is to restrict the plasma flow in the transverse direction, which points to the fact that condensation always occurs along the magnetic field lines as the parallel velocity is mostly unrestricted. However, even a small transverse velocity component can make a considerable difference in the final condensation state (Hennebelle & Pérault, 1999). As a result, we can expect that the thermal instability will be stifled due to the flow restrictions imposed by the magnetic field. However, as the total energy content must be conserved, we certainly expect a field aligned, more diffused condensation state and the resultant velocity field will not be highly ordered leading to a turbulent-like plasma flow, as can be seen from the right figure of the second row of Fig.7 (see next Subsection).
We now consider the case for a ‘very strong’ magnetic field with $\beta=0.1$, and show the evolution of density after $1500\,\rm{Myr}$ in the presence of thermal conduction with $a_{p}=3$ (first row panels) as well as in the absence of thermal conduction for three different values of pressure anisotropy parameter, $a_{p}=0,1,\,\rm{and}\,3$ (rest of the panels) in Fig.6. For each left panel, the corresponding right panel presents a zoomed-in view of a selected region from the density map. In the presence of conduction with a ‘very strong’ field, the nature of initial evolution of density remains the same as in the previous cases with condensations largely elongated in the direction parallel to the magnetic field and a slower start of the growth of instability. However, as it approaches the nonlinear saturation stage, some of the condensations start getting elongated in the direction perpendicular to the field, which is shown by a zoomed in portion in Fig.6(b). In contrast to this, in the absence of thermal conduction, condensations are observed to be very different. These filaments are comparatively shorter, less diffused and get elongated mostly in the perpendicular direction of the field as shown in Fig.6(c) and Fig.6(d). Similar nature of the condensations in the absence of thermal conduction was earlier reported by Wareing et al. (2016). The inclusion of anisotropic thermal conduction in the model predominantly channels the formation of condensations along the direction of background magnetic field and suppresses growth along the direction perpendicular to the field (Choi & Stone, 2012).
As we understand now, in absence of thermal conduction, there is no mechanism to erase temperature variations. In this case, apparently a strong magnetic field greatly reduces motion of the plasma across field lines, allowing density perturbations along the field lines which can cause sharp density gradient along the field lines resulting filamentary structures which appear to form across field lines (Wareing et al., 2016). In contrast to these, a large parallel thermal conductivity damps temperature perturbations along the field lines. So, small $k_{\parallel}$ (large parallel wavelength) modes will be less stabilised and will grow fast, which results in field-aligned structures. As we have seen, in both cases our simulation produces the desired results.
At this point, we would like to also emphasise on the effect of pressure anisotropy on the condensations, which is more pronounced for a strong magnetic field. We recall that in our case, the pressure anisotropy can be modelled with a single parameter $a_{p}$ – while $a_{p}=0$ corresponds to $p_{\parallel}=p_{\perp}$, a positive $a_{p}$ corresponds to stronger $p_{\parallel}$. Let us consider the panels (d, f, h) of Fig.6 (right column) which show the corresponding zoomed-in portion of the density filaments with zero thermal conductivity for $a_{p}=3,1$, and 0 (top to bottom). What we see is that as pressure anisotropy becomes stronger (larger $p_{\parallel}$), the filamentations become longer, more connected and get elongated across the magnetic field line, while for lower anisotropy, we have scattered and more or less isotropic filamentation. This is a stark deviation from earlier observations (Wareing et al., 2016), for lower $a_{p}$. Going by our earlier argument, we see that stronger parallel pressure suppresses compressions along the field lines.